mass = 9.9876±0.0004 g mass = 9.9866±0.0004 g mass = 9.9873±0.0004 g mass = 9.9878±0.0004 g...

  1. mass = 9.9876±0.0004 g
  2. mass = 9.9866±0.0004 g
  3. mass = 9.9873±0.0004 g
  4. mass = 9.9878±0.0004 g
  5. mass = 9.9864±0.0004 g

A student was calibrating a 10-mL pipette which was guaranteed by manufacturers to deliver 10.00 ± 0.02 mL if used correctly. He measured the mass of water delivered by the pipette five times and obtained the masses shown on the right. The mean mass is 9.98714 g.

a What is the absolute uncertainty associated with this mean mass? (use 1/2 range)

b He used a conversion factor (1 g = 1.0021 mL) to convert mass to volume.
What is the mean volume?
Answer #1

SOLO m085 = 9.9876 + 0.0004 g mass = 9.9866 0.00048 ma58 19.9873 $ 0.0004 g ma88 - 9.9878 + 0.00048 maBB 9.9864 + 0.0004 8. CSteps («; -ā), izl S2 } (xi - x)? N-1 S = 3.78 S 0.000 6148 Thus, uncertainty associated with this mean magg 0.00062 g He use